Leak Inc. forecasts the free cash flows (in millions) shown below. If the weighted average cost of capital is 11% and FCF is expected to grow at a rate of 5% after Year 2, what is the Year 0 value of operations, in millions? Assume that the ROIC is expected to remain constant in Year 2 and beyond (and do not make any half-year adjustments).
Year: 1 2
Free cash flow: -$50 $100
a. $1,456
b. $1,529
c. $1,606
d. $1,686
e. $1,770
1
Expert's answer
2012-08-17T07:49:04-0400
The answer is: e ($ 1,770)




Explanation:

Terminal alue in Year 2 = CF3/(Ke-g), where

CF3=cash flow in year 3=100x1,05=105

Ke=required return=11%

g=growth rate=5%

Terminal Value in Year 2 = 105/(11%-5%) = 1,770